# Data-adaptive structural change-point detection via isolation

### v2

@@ 7.2 Euro to British pound exchange rate
The DAIS algorithm for changes in the slope is applied to the Euro to British pound exchange rate. The data are the weekly close prices adjusted for splits for the period between 07/04/2014 and 12/09/2024. The results can be seen in Figure 4. The first plot is the observed signal, while the other plots are the estimated piecewise-linear signals from DAIS, ID and CPOP. The change-points which are discussed in the next paragraph are indicated by dashed lines in the plot of the observed signal.
@@
Figure 4: The observed time series data of Euro to British pound sterling exchange rate and fitted piecewise-linear signal obtained by applying DAIS (red), ID (green) and CPOP (blue).
@@ 8 Conclusions
In this project, DAIS is introduced, which is a new data-adaptive method for detecting structural changes in a given data sequence. The first step of the method is to identify the location of the largest difference in the sequence, as defined in (2) for the cases of changes in the mean and changes in the slope, in an effort to start searching from an area around the most apparent change-point in the case that there is one. This step is what makes our method data-adaptive and differentiates it from the competitors. Expanding intervals around the location of the largest difference, in absolute value, are used to achieve isolation of the change-points, which enhances the detection power. Theoretical results regarding the consistency of the number of change-points detected and the accuracy of their estimated locations are provided. The simulation results presented indicate that DAIS is at least as accurate as the state-of-the-art competitors, with its data-adaptive nature being advantageous regarding the detection of the true change-points in difficult structures, for example when the change-points are really close to each other. DAIS uses intervals around the location of the largest difference and so the change-points are, with high probability, near the midpoint of these intervals, which results in the power of the contrast function being maximized. For the proof of the theoretical properties of DAIS we require that the location of the largest difference is at a location where detection, while the change-point is isolated in an interval, is guaranteed. However, for the correct choice of the parameters, the points that are problematic are just the midpoints between two consecutive change-points. In practice, the algorithm performs at least as well as other widely-used algorithms in the literature, exhibiting uniformly good behaviour across different signal structures. Finally, some extensions to more difficult signal structures are discussed, which include heavy tailed noise, temporal dependence and the extension of the method to multivariate signals.
@@ Appendix A Signals used in the simulation study
The signals used in the simulations of Section 5 are reported below. For the case of piecewise-constant the signals $f_{t}$ in (1) are the following:
@@
    $many\_cpts\_mix$: piecewise-constant signal of length 75 with 11 change-points at 5, 12, 17, 25, 31, 38, 44, 50, 56, 61, 67 with values between the change-points 0, 5, 0, 6, 0, 4, 0, 5, 0, 6, 0, 4. The standard deviation is $\sigma=1$.
@@
The signals used for the case of piecewise-linear $f_{t}$ are:
@@
    $wave1$: continuous piecewise-linear signal of length 1408 with 7 change-points at 256, 512, 768, 1024, 1152, 1280, 1344 with changes in slope $-1/64$, $2/64$, $-3/64$, $4/64$, $-5/64$, $6/64$, $-7/64$. The starting intercept is $f_{1}=1$ and slope $f_{2}-f_{1}=1/256$. The standard deviation is $\sigma=1$.
@@
    $wave2$: continuous piecewise-linear signal of length 1500 with 99 change-points at 15, 30, $\ldots$, 1485 with changes in slope -1, 1, -1 $\ldots$, -1. The starting intercept is $f_{1}=-1/2$ and slope $f_{2}-f_{1}=1/40$. The standard deviation is $\sigma=1$.
@@
    $wave3$: continuous piecewise-linear signal of length 840 with 119 change-points at 7, 14, 21, $\ldots$, 833 with changes in slope -1, 1, -1 $\ldots$, -1. The starting intercept is $f_{1}=-1/2$ and slope $f_{2}-f_{1}=1/32$. The standard deviation is $\sigma=0.3$.
Appendix B Further simulations
Some further simulation results are presented in Tables 8 and 9 for piecewise-constant signals and Tables 10 and 11 for continuous piecewise-linear signals. The signals used are:
  • (S11)
  • (S12)
  • (S13)
  • (S14)
  • (S15)
  • (S16)
    $wave4$: continuous piecewise-linear signal of length 200 with 9 change-points at 20, 40, $\ldots$, 180 with changes in slope 1/6, 1/2, -3/4, -1/3, -2/3, 1, 1/4, 3/4, -5/4. The starting intercept is $f_{1}=-1$ and slope $f_{2}-f_{1}=1/32$. The standard deviation is $\sigma=0.3$.
  • (S17)
    $wave5$: continuous piecewise-linear signal of length 350 with 50 change-points at 7, 14, $\ldots$, 343 with changes in slope -2.5, 2.5, $\ldots$, 2.5, -2.5. The starting intercept is $f_{1}=-0$ and slope $f_{2}-f_{1}=1$. The standard deviation is $\sigma=1$.
The general conclusions form the tables are that DAIS performs at least as well as the competitors, often having an advantage in computational time and accuracy compared to the best performing competitors. Signal (S11), in Table 8 has no change-points. Most algorithms have very good results with negligible differences in the accuracy in the number of change-points detected, WBS_th with the threshold constant $C=\sqrt{2}$ seems to overestimate. Similar conclusions hold for signal (S12), in which the big length of the signal forces some algorithms, especially MSCP, to be extremely slow. Moreover, comparing DAIS with ID_th which use the same value for the expansion parameter $\lambda_{T}$, it is evident that DAIS requires a smaller number of expansions and thus is faster. Signal (S13) in Table 9 indicates that in this difficult structure, where pairs of change-points are close and move in opposite directions such that they ‘cancel’ each-other out, DAIS and MOSUM have the best performance, in terms of accuracy in the number of change-points detected and MSE. For the signal (S14) all algorithms perform similarly, with DAIS being one of the fastest methods. The low computational complexity of DAIS can also be seen in Table 11.
Table 8: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signals (S11) and (S12). The average MSE and computational times are also given.
$\hat{N}-N$
Method Signal -1 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 99 0 1 0 $1.87\times 10^{-4}$ 0.237
ID_th 0 84 8 7 1 $5.26\times 10^{-4}$ 0.450
ID_ic 0 99 1 0 0 $1.68\times 10^{-4}$ 0.080
WBS_th 0 52 26 20 2 $7.00\times 10^{-4}$ 0.067
WBS_ic 0 100 0 0 0 $1.39\times 10^{-4}$ 0.066
WBS2 (S11) 0 91 5 3 1 $2.75\times 10^{-4}$ 3.51
PELT 0 100 0 0 0 $1.39\times 10^{-4}$ 0.003
NOT 0 100 0 0 0 $1.39\times 10^{-4}$ 0.066
MOSUM 0 94 2 4 0 $2.96\times 10^{-4}$ 0.028
MSCP 0 100 0 0 0 $1.39\times 10^{-4}$ 83.25
SeedBS_th 0 12 23 25 40 $12.9\times 10^{-4}$ 0.027
SeedBS_ic 0 99 1 0 0 $1.68\times 10^{-4}$ 0.027
DP_univar 0 0 0 0 100 $372\times 10^{-4}$ 30.70
DAIS 0 99 0 1 0 $4.33\times 10^{-4}$ $43.10\times 10^{-4}$ 0.322
ID_th 0 88 1 10 1 $7.77\times 10^{-4}$ $35.80\times 10^{-4}$ 0.416
ID_ic 0 100 0 0 0 $4.25\times 10^{-4}$ $1.09\times 10^{-4}$ 0.110
WBS_th 0 75 19 6 0 $5.15\times 10^{-4}$ $626.00\times 10^{-4}$ 0.107
WBS_ic 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.107
WBS2 (S12) 0 92 3 4 1 $5.46\times 10^{-4}$ $17.60\times 10^{-4}$ 5.950
PELT 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.004
NOT 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.103
MOSUM 0 92 2 6 0 $5.49\times 10^{-4}$ $203.60\times 10^{-4}$ 0.049
MSCP 6 93 1 0 0 $341.35\times 10^{-4}$ $334.74\times 10^{-4}$ 276.294
SeedBS_th 0 10 23 35 42 $10.2\times 10^{-4}$ $3140\times 10^{-4}$ 0.0418
SeedBS_ic 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.0418
DP_univar 0 0 0 0 100 $368\times 10^{-4}$ $4840\times 10^{-4}$ 190.80
$\hat{N}-N$
Method Signal $\leq-3$ $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 15 3 80 0 2 0 0.0321 0.067 0.004
ID_th 0 2 2 75 8 7 6 0.0323 0.037 0.003
ID_ic 2 34 1 62 1 0 0 0.0350 0.143 0.002
WBS_th 0 0 1 31 28 26 14 0.0336 0.088 0.025
WBS_ic 2 33 1 64 0 0 0 0.0337 0.140 0.025
WBS2 (S13) 0 6 5 70 9 6 4 0.0306 0.132 0.479
PELT 26 65 0 9 0 0 0 0.0596 0.437 0.001
NOT 2 30 1 67 0 0 0 0.0333 0.128 0.024
MOSUM 0 13 2 79 3 3 0 0.0287 0.063 0.015
MSCP 95 5 0 0 0 0 0 0.1390 0.593 2.610
SeedBS_th 0 0 1 18 27 14 40 0.0381 0.097 0.013
SeedBS_ic 1 29 2 63 3 1 1 0.0331 0.122 0.013
DP_univar 0 0 0 2 6 4 88 0.0602 0.126 0.139
DAIS 0 0 1 94 4 1 0 2.72 ×$\times 10^{-2}$ 0.010 0.002
ID_th 0 0 1 84 12 2 1 2.72 ×$\times 10^{-2}$ 0.011 0.001
ID_ic 0 0 0 87 12 1 0 2.46 ×$\times 10^{-2}$ 0.009 0.002
WBS_th 0 0 0 56 21 19 4 2.58 ×$\times 10^{-2}$ 0.016 0.008
WBS_ic 0 0 0 93 7 0 0 2.35 ×$\times 10^{-2}$ 0.008 0.008
WBS2 (S14) 0 0 1 96 3 0 0 2.31 ×$\times 10^{-2}$ 0.008 0.088
PELT 1 6 3 90 0 0 0 2.63 ×$\times 10^{-2}$ 0.016 0.001
NOT 0 0 0 95 5 0 0 2.47 ×$\times 10^{-2}$ 0.009 0.021
MOSUM 0 0 0 95 3 2 0 2.26 ×$\times 10^{-2}$ 0.008 0.008
MSCP 100 0 0 0 0 0 0 21.10 ×$\times 10^{-2}$ 0.261 0.242
SeedBS_th 0 0 1 49 25 15 10 2.68 ×$\times 10^{-2}$ 0.017 0.011
SeedBS_ic 0 0 0 93 5 2 0 2.40 ×$\times 10^{-2}$ 0.090 0.012
DP_univar 49 18 2 31 0 0 0 8.89 ×$\times 10^{-2}$ 0.210 0.003
Table 10: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signal (S15). The average MSE and computational times are also given.
$\hat{N}-N$
Method Signal 0 1 2 $\geq 3$ MSE Time (s)
DAIS 100 0 0 0 $2.12\times 10^{-3}$ 0.032
ID_th 83 10 4 3 $2.45\times 10^{-3}$ 0.066
ID_ic 100 0 0 0 $2.12\times 10^{-3}$ 0.007
CPOP 100 0 0 0 $2.45\times 10^{-3}$ 3.220
NOT (S15) 100 0 0 0 $2.31\times 10^{-3}$ 0.062
MARS 0 100 0 0 $5.25\times 10^{-3}$ 0.006
TF 7 44 37 12 $5.29\times 10^{-3}$ 0.429
TS 100 0 0 0 $2.12\times 10^{-3}$ 0.618
local_poly 10 5 17 68 $6.47\times 10^{-3}$ 33.000
$\hat{N}-N$
Method Signal $\leq-15$ $(-15,-2]$ $-1$ 0 1 $[2,15)$ $\geq 15$ MSE $d_{H}$ Time (s)
DAIS 0 0 0 96 4 0 0 0.022 0.104 0.003
ID_th 0 0 0 76 19 5 0 0.021 0.149 0.002
ID_ic 0 0 0 99 1 0 0 0.013 0.097 0.004
CPOP 0 0 0 87 12 1 0 0.081 0.122 0.017
NOT (S16) 0 0 1 96 1 2 0 0.015 0.097 0.082
MARS 0 84 13 3 0 0 0 3.720 2.160 0.003
TF 0 0 0 0 0 21 79 0.024 0.433 0.097
TS 0 0 0 96 4 0 0 0.094 0.198 0.110
local_poly 0 100 0 0 0 0 0 0.331 0.956 0.095
DAIS 0 1 2 96 1 0 0 0.461 0.300 0.012
ID_th 0 0 0 92 8 0 0 0.482 0.294 0.006
ID_ic 39 1 15 44 1 0 0 3.100 19.50 0.025
CPOP 0 0 0 89 11 0 0 1.650 0.323 0.025
NOT (S17) 100 0 0 0 0 0 0 6.650 49.00 0.039
MARS 100 0 0 0 0 0 0 6.590 48.00 0.003
TF 0 0 0 0 0 82 18 1.200 0.323 0.110
TS 100 0 0 0 0 0 0 6.650 49.00 0.197
local_poly 100 0 0 0 0 0 0 6.610 5.220 0.639
Appendix C Proof of Theorem 1
Before proving Theorem 1, we introduce some more notation. For convinience, we denote (10) by $\tilde{X}^{b}_{s,e}=\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum^{b}_{t=s}X_{t}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum^{e}_{t=b+1}X_{t}.$ Similarly, we denote the CUSUM of the unobserved signal as $\tilde{f}^{b}_{s,e}=\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum^{b}_{t=s}f_{t}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum^{e}_{t=b+1}f_{t}$. For the following proof, the contrast vector $\boldsymbol{\psi_{s,e}^{b}}=\left(\psi_{s,e}^{b}(1),\psi_{s,e}^{b}(2),\dots,\psi_{s,e}^{b}(T_{1})\right)$ is defined through the contrast function
@@ Appendix C Proof of Theorem 1
$B_{T}=\left\{\max_{1\leq s\leq b<e<{T}}\max_{\begin{subarray}{c}r_{j-1}<s\leq r_{j}\\ r_{j}<e\leq r_{j+1}\\ s\leq b<e\end{subarray}}\frac{\left|\left\langle\boldsymbol{\psi_{s,e}^{b}}\langle\boldsymbol{f},\boldsymbol{\psi_{s,e}^{b}}\rangle-\boldsymbol{\psi_{s,e}^{r}}\langle\boldsymbol{f},\boldsymbol{\psi_{s,e}^{r}}\rangle,\boldsymbol{\epsilon}\right\rangle\right|}{\|\boldsymbol{\psi_{s,e}^{b}}\langle\boldsymbol{f},\boldsymbol{\psi_{s,e}^{b}}\rangle-\boldsymbol{\psi_{s,e}^{r}}\langle\boldsymbol{f},\boldsymbol{\psi_{s,e}^{r}}\rangle\|_{2}}\leq\sqrt{8\log T}\right\}.$ (30)
@@
where $\underline{C}$ satisfies assumption (A2), $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}\sqrt{\log T}$ and $3/2\leq n\leq\delta_{T}/2\lambda_{T}$ as defined in Section 2.3. For $j\in\{1,2,\ldots,N\}$ define $I^{L}_{j}$ and $I^{R}_{j}$ as in (5). The location of the largest difference in an interval $[s,e]$, $1\leq s<e\leq T$, is defined as $d_{s,e}=\textrm{argmax}_{t\in\{s,s+1,\dots,e-1\}}\{\lvert X_{t+1}-X_{t}\rvert\}$ as in (2) and for $K^{l}=\lceil\frac{d_{s,e}-s+1}{\lambda_{T}}\rceil$, $K^{r}=\lceil\frac{e-d_{s,e}+1}{\lambda_{T}}\rceil$, $K^{\max}=\max\{K^{l},K^{r}\}$, define
@@
Since the length of the intervals in (5) is $\delta_{T}/2n$ and $\lambda_{T}\leq\delta_{T}/2n$, for $m,k\in\{0,1,\ldots,K^{\max}\}$ we ensure that there exists at least one $m$ and at least one $k$ such that $c_{m}^{l}\in I^{L}_{j}$ and $c_{k}^{r}\in I^{R}_{j}$ for all $j\in\{1,2,\ldots,N\}$ and for all possible locations of the largest difference.
At the beginning of DAIS, $s=1,e=T$ and the first change-point that will get detected depends on the value of $d_{1,T}$. As already explained in Section 2.3, for $j\in\{1,\ldots,N-1\}$, the largest difference $d_{s,e}$ will be at most at a distance $\delta_{1,T}^{j}$ from the nearest change-point $r_{j}$ or $r_{j+1}$, where $\delta_{1,T}^{j}\leq\frac{\tilde{\delta}_{j}}{2}-\frac{3\delta_{T}}{4n}$ for $\tilde{\delta}_{j}=r_{j+1}-r_{j}$. The first point to get detected will be the point that is closest to the largest difference $d_{1,T}$. The interval where the detection of this change-point occurs, cannot contain more than one change-points, as was explained in Section 2.3 and is proved at (C).
We will show that there exists an interval $[c_{m}^{l},c_{k}^{r}]$, for $m,k\in\{0,1,\ldots,K^{\max}\}$, such that the first change-point to get detected, $r_{J}$, is isolated, assuming that $d_{1,T}>r_{J}$. The same approach works for $d_{1,T}\leq r_{J}$. Without loss of generality, we suppose that $J\in\left\{1,2,\ldots,N-1\right\}$. The result in the case that $J=N$ is easier to prove and so it is skipped. Note that since the closest change-point to the largest difference is $r_{J}$, it must hold that $d_{1,T}^{J}=d_{1,T}-r_{J}$. We are now considering the largest possible interval that may be checked before detection occurs in the sense that the largest number of expansions will be performed. Detection is guaranteed when checking this interval, but it may occur in any interval smaller than this. Showing that the change-point will be isolated in the largest interval means that it will always be isolated. Since we are considering $d_{1,T}>r_{J}$, the worst case scenario occurs for $m=k$, $c_{m}^{l}\in I^{L}_{J}$ and $c_{k}^{r}-r_{J}>\delta_{T}/2n$, but it doesn’t necessarily hold that $c_{k}^{r}\in I^{R}_{J}$. So, $r_{J}-c_{m}^{l}\leq\delta_{T}/n$ and $c_{k}^{r}-d_{1,T}+1=d_{1,T}-c_{m}^{l}$. It follows that
@@
$<2\Bigl{(}\delta_{1,T}^{J}+\frac{\delta_{T}}{n}\Bigr{)}\leq\tilde{\delta}_{J}+\frac{\delta_{T}}{2n}.$ (33)
Now, since $r_{J}-c_{m}^{l}\leq\delta_{T}/n$, there can be at most one change-point in the interval $[c_{m}^{l},r_{J}]$. Considering the interval $[r_{J},c_{k}^{r}]$, using (C) and $r_{J}-c_{m}^{l}>\delta_{T}/2n$, it holds that $c_{k}^{r}-r_{J}=c_{k}^{r}-c_{m}^{l}+c_{m}^{l}-r_{J}<\tilde{\delta}_{J}$, so $c_{k}^{r}<r_{J+1}$ and $r_{J}$ is isolated in $[r_{J},c_{k}^{r}]$. The two results combined prove that the change-point is isolated in the interval $[c_{m}^{l},c_{k}^{r}]$.
@@
then using assumption (A2) and the results in (34) (35), (37), we have that
@@
Without loss of generality, assume that $b^{\ast}\in[r_{J},c_{k^{\ast}}^{r})$ and a similar approach holds when $b^{\ast}\in[c_{m^{\ast}}^{l},r_{J})$. Using Lemma 1, we have that for the left-hand side of (C),
@@
Using (41), (42) and (C), we can conclude that (C) is satisfied if $\Lambda>8\log T+\sqrt{2}\sqrt{\Lambda}\sqrt{8\log T}$ is satisfied, which has solution

### v3

@@ 7.2 Euro to British pound exchange rate
The DAIS algorithm for changes in the slope is applied to the Euro to British pound exchange rate. The data are the weekly close prices adjusted for splits for the period between 07/04/2014 and 12/09/2024, obtained from https://finance.yahoo.com/quote/EURGBP%3DX/. The results can be seen in Figure 4. The first plot is the observed signal, while the other plots are the estimated piecewise-linear signals from DAIS, ID and CPOP. The change-points which are discussed in the next paragraph are indicated by dashed lines in the plot of the observed signal.
@@
Figure 4: The observed time series data of Euro to British pound sterling exchange rate and fitted piecewise-linear signal obtained by applying DAIS (red), ID (green) and CPOP (blue). The x-axis labels correspond to the year.
@@ 8 Conclusions
In this paper, DAIS is introduced, which is a new data-adaptive method for detecting structural changes in a given data sequence. The first step of the method is to identify the location of the largest difference in the sequence, as defined in (2) for the cases of changes in the mean and changes in the slope, in an effort to start searching from an area around the most apparent change-point in the case that there is one. This step is what makes our method data-adaptive and differentiates it from the competitors. Expanding intervals around the location of the largest difference, in absolute value, are used to achieve isolation of the change-points, which enhances the detection power. Theoretical results regarding the consistency of the number of change-points detected and the accuracy of their estimated locations are provided. The simulation results presented indicate that DAIS is at least as accurate as the state-of-the-art competitors, exhibiting uniformly good behavior across different signal structures. Its data-adaptive nature is advantageous regarding the detection of the true change-points in difficult structures. DAIS uses intervals around the location of the largest difference and so the change-points are, with high probability, near the midpoint of these intervals, which results in the power of the contrast function being maximized. Such an interval expansion approach around the location of the true change-points, offers also an advantage in terms of the computational complexity of the algorithm, especially in cases with few change-points. For the proof of the theoretical properties of DAIS we require that the location of the largest difference is at a location where detection, while the change-point is isolated in an interval, is guaranteed. However, in practice, the change-points can still potentially get detected even when this assumption fails. Detection could occur in an interval in which the change-point is isolated, but whose end-points are not far enough from the true location of the change-point to guarantee detection. It could also occur in an interval in which the change-point is not isolated. In such intervals, as described in both cases, the detection power is reduced, but detection can still occur.
Finally, we consider some extensions of DAIS to more difficult structures. This includes heavy tailed noise and temporal dependence. The case of multivariate signals is also discussed and further analyzed, with the algorithm MDAIS being proposed. Under some assumptions that are equivalent to the ones for the univariate signals, a consistency result is provided for the case of possible changes in the mean vector of an underlying $d-$dimensional signal.
@@ Appendix A Signals used in the simulation study
The signals used in the simulations of Section 5 are reported below. For the case of changes in the mean (piecewise-constancy), the signals $f_{t}$ in (1) are the following:
@@
  • (S2)
    $small\_dist2$ piecewise-constant signal of length 135 with 2 change-points at 30, 35 with values between change-points 0, 2.3, 8. The standard deviation is $\sigma=1$.
@@
    $mix2$: piecewise-constant signal of length 75 with 11 change-points at 5, 12, 17, 25, 31, 38, 44, 50, 56, 61, 67 with values between the change-points 0, 5, 0, 6, 0, 4, 0, 5, 0, 6, 0, 4. The standard deviation is $\sigma=1$.
@@
Under the case of continuous, piecewise-linearity, the signals $f_{t}$ used are:
@@
    $wave1$: continuous, piecewise-linear signal of length 1408 with 7 change-points at 256, 512, 768, 1024, 1152, 1280, 1344 with changes in slope $-1/64$, $2/64$, $-3/64$, $4/64$, $-5/64$, $6/64$, $-7/64$. The starting intercept is $f_{1}=1$ and slope $f_{2}-f_{1}=1/256$. The standard deviation is $\sigma=1$.
@@
    $wave2$: continuous, piecewise-linear signal of length 1500 with 99 change-points at 15, 30, $\ldots$, 1485 with changes in slope -1, 1, -1 $\ldots$, -1. The starting intercept is $f_{1}=-1/2$ and slope $f_{2}-f_{1}=1/40$. The standard deviation is $\sigma=1$.
@@
    $wave3$: continuous, piecewise-linear signal of length 840 with 119 change-points at 7, 14, 21, $\ldots$, 833 with changes in slope -1, 1, -1 $\ldots$, -1. The starting intercept is $f_{1}=-1/2$ and slope $f_{2}-f_{1}=1/32$. The standard deviation is $\sigma=0.3$.
Appendix B Proof of Theorem 1
Before proving Theorem 1, we introduce some more notation. For convenience, we denote (15) by $\tilde{X}^{b}_{s,e}=\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum^{b}_{t=s}X_{t}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum^{e}_{t=b+1}X_{t}.$ Similarly, we denote the CUSUM of the unobserved signal as $\tilde{f}^{b}_{s,e}=\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum^{b}_{t=s}f_{t}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum^{e}_{t=b+1}f_{t}$. We also work under (12), (13) and (14). For the following proof, the contrast vector $\boldsymbol{\psi_{s,e}^{b}}=\left(\psi_{s,e}^{b}(1),\psi_{s,e}^{b}(2),\dots,\psi_{s,e}^{b}(T_{1})\right)$ is defined through the contrast function
@@ Appendix B Proof of Theorem 1
$B_{T}=\left\{\max_{j=1,2,\ldots,N}\max_{\begin{subarray}{c}r_{j-1}<s\leq r_{j}\\ r_{j}<e\leq r_{j+1}\\ s\leq b<e\end{subarray}}\frac{\left|\left\langle\boldsymbol{\psi_{s,e}^{b}}\langle\boldsymbol{f},\boldsymbol{\psi_{s,e}^{b}}\rangle-\boldsymbol{\psi_{s,e}^{r}}\langle\boldsymbol{f},\boldsymbol{\psi_{s,e}^{r}}\rangle,\boldsymbol{\epsilon}\right\rangle\right|}{\|\boldsymbol{\psi_{s,e}^{b}}\langle\boldsymbol{f},\boldsymbol{\psi_{s,e}^{b}}\rangle-\boldsymbol{\psi_{s,e}^{r}}\langle\boldsymbol{f},\boldsymbol{\psi_{s,e}^{r}}\rangle\|_{2}}\leq\sqrt{8\log T}\right\}.$ (36)
@@
where $\underline{C}$ satisfies Assumption (A2), $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}\sqrt{\log T}$ and $n\geq 3/2$ as defined in Section 2.3. For $j\in\{1,2,\ldots,N\}$ define $I^{L}_{j}$ and $I^{R}_{j}$ as in (5). The location of the largest difference in an interval $[s,e]$, $1\leq s<e\leq T$, is defined as $d_{s,e}=\textrm{argmax}_{t\in\{s,s+1,\dots,e-1\}}\{\lvert X_{t+1}-X_{t}\rvert\}$ as in (2) and for $K^{l}=\lceil\frac{d_{s,e}-s+1}{\lambda_{T}}\rceil$, $K^{r}=\lceil\frac{e-d_{s,e}+1}{\lambda_{T}}\rceil$, $K^{\max}=\max\{K^{l},K^{r}\}$, define
@@
Since the length of the intervals in (5) is $\delta_{T}/2n$, $\lambda_{T}\leq\delta_{T}/2n$ ensures that there exists at least one $m\in\{0,1,\ldots,K^{\max}\}$ and at least one $k\in\{1,2\ldots,K^{\max}\}$ such that $c_{m}^{l}\in I^{L}_{j}$ and $c_{k}^{r}\in I^{R}_{j}$ for all $j\in\{1,2,\ldots,N\}$ and for all possible locations of the largest difference.
At the beginning of DAIS, $s=1,e=T$ and the first change-point that will get detected depends on the value of $d_{1,T}$. As already explained in Section 2.3, for $j\in\{1,\ldots,N-1\}$, the largest difference $d_{s,e}$ will be at most at a distance $\delta_{1,T}^{j}$ from the nearest change-point $r_{j}$ or $r_{j+1}$, where $\delta_{1,T}^{j}\leq\frac{\tilde{\delta}_{j}}{2}-\frac{3\delta_{T}}{4n}$ for $\tilde{\delta}_{j}=r_{j+1}-r_{j}$. The first point to get detected will be the point that is closest to the largest difference $d_{1,T}$. The interval where the detection of this change-point occurs, cannot contain more than one change-points, as was explained in Section 2.3 and is proved at (B).
We will show that there exists an interval $[c_{m}^{l},c_{k}^{r}]$, for $m\in\{0,1,\ldots,K^{\max}\}$ and $k\in\{1,2\ldots,K^{\max}\}$, such that the first change-point to get detected, $r_{J}$, is isolated, assuming that $d_{1,T}>r_{J}$. A similar approach works for $d_{1,T}\leq r_{J}$. The differences arise from the fact that the expansions are always performed starting from the right, which implies that if the change-point is to the right of $d_{1,T}$, an uneven number of expansions will be performed before the end-point of the interval lies in (5). Without loss of generality, we suppose that $J\in\left\{1,2,\ldots,N-1\right\}$. The result in the case that $J=N$ is easier to prove and so it is skipped, since, as can be seen by (9), the change-point will certainly be isolated when detection occurs if $d_{1,T}>r_{N}$. Note that since the closest change-point to the largest difference is $r_{J}$, it must hold that $\delta_{1,T}^{J}=d_{1,T}-r_{J}$. We are now considering the largest possible interval that may be checked before detection occurs in the sense that the largest number of expansions will be performed. Detection is guaranteed when checking this interval, but it may occur in any interval smaller than this. Showing that the change-point will be isolated in the largest interval means that it will also be isolated in any of the smaller intervals. Since we are considering $d_{1,T}>r_{J}$, the worst case scenario, (meaning when DAIS checks the largest interval) occurs for $m=k$, $c_{m}^{l}\in I^{L}_{J}$ and $c_{k}^{r}-r_{J}>\delta_{T}/2n$, but it does not necessarily hold that $c_{k}^{r}\in I^{R}_{J}$. So, $r_{J}-c_{m}^{l}\leq\delta_{T}/n$ and $c_{k}^{r}-d_{1,T}+1=d_{1,T}-c_{m}^{l}$. It follows that
@@
$<2\Bigl{(}\delta_{1,T}^{J}+\frac{\delta_{T}}{n}\Bigr{)}\leq\tilde{\delta}_{J}+\frac{\delta_{T}}{2n},$ (39)
where we used (6) in the last step. Now, since $r_{J}-c_{m}^{l}\leq\delta_{T}/n$, there can be at most one change-point in the interval $[c_{m}^{l},r_{J}]$. Considering the interval $[r_{J},c_{k}^{r}]$, using (B) and $r_{J}-c_{m}^{l}>\delta_{T}/2n$, it holds that $c_{k}^{r}-r_{J}=c_{k}^{r}-c_{m}^{l}+c_{m}^{l}-r_{J}<\tilde{\delta}_{J}$, so $c_{k}^{r}<r_{J+1}$ and $r_{J}$ is isolated in $[r_{J},c_{k}^{r}]$. The two results combined prove that the change-point is isolated in the interval $[c_{m}^{l},c_{k}^{r}]$.
@@
then, using Assumption (A2) and the results in (40) (41), (43), we have that
@@
Without loss of generality, assume that $b^{\ast}\in[r_{J},c_{k^{\ast}}^{r})$ and a similar approach holds when $b^{\ast}\in[c_{m^{\ast}}^{l},r_{J})$. Using Lemma 1, we have that for the left-hand side of (B),
@@
Using (47), (48) and (B), we can conclude that (B) is satisfied if $\Lambda>8\log T+\sqrt{2}\sqrt{\Lambda}\sqrt{8\log T}$ is satisfied, which has solution